Use the order of operations to simplify: 2.35 – (3.09 – 1.09)² ÷ 0.5.
I am coming up with -3.3 as the answer. Is this correct. Please verify steps. Thanks.

Answer by mbmodak(32)

2.35 – ((3.09 – 1.09)^²/ 0.5))
(3.09-1.09)^2=(2)^2=4
4/05=8
2.35-8=-5.65